IPP Offline Activities

Have a long list of offline tasks

1. Management of CDF Offline group (PKS - 0.5 FTE)

2. CDFSIM (Sganos)

3. Geometry (Ragan)

4. Bank dump code - BNK (Sganos)

5. Physics Utilities - PU and UTI (Warburton)

6. DF (Warburton)

7. Generator area (PKS)

Have to reexamine list of responsibilities for Run II

• Can’t be stretched too thin

• Can’t just drop these either

One Proposed Plan

List of tasks is probably too large at the moment

• Have to prioritise tasks

• identify which ones are already orphans – CDFSIM

– BNK

– Generator (maybe)

• Which ones are one their way...

– DF and Utilities?

Address our offline involvement for Run II

• Have to redesign offline system

• No more VMS

• Integrate non-FORTRAN/77 languages